n educator wants to eve uste uur o erent rne uds arned at recuoriq classroom. The results are given in the table. t reducing the me children ฯpend ch uff as s in he oesroom. Iu test es methods she irmplements une rme ายป r18dch o our ฯn ar classrooms and recor s the mie spent u task in mimu each 4 3 0 4 ) Oomplele the F-Lable, (Ruu yuur ariswers to two detiral places oif Betwee groups (arrar
Answer #1

Ans. The given problem is studied under the Analysis of variance(ANOVA), i.e., we need to calculate the variation between and within the groups.

A

(x_{1})

mathbf{A}

x_{1}^{2}

mathbf{B}

T2)

x_{2}^{2}

mathbf{C}

T3

x_{3}^{2}

mathbf{D}

(x_{4})

x_{4}^{2}

3

9 4 16 4 16 4 16
1 1 1 1 4 16 2 4
4 16 3 9 0 0 4 16
2 4 2 4 5 25 4 16
0 0 5 25 6 36 5 25
5 25 0 0 4 16 3 9
3 9 0 0 3 9 3 9
5 25 4 16 0 0 2 4
1 1 1 1 5 25 2 4
Sigma x_{1}=24 mathbf{Sigma x_{1}^{2}=90} Sigma x_{2}=20 mathbf{Sigma x_{2}^{2}=72} Sigma x_{3}=31 mathbf{Sigma x_{3}^{2}=143} Sigma x_{4}=29 mathbf{Sigma x_{4}^{2}=103}
mathbf{ar{x_{1}}=2.6667} mathbf{ar{x_{2}}=2.2222} mathbf{ar{x_{3}}=3.4444} mathbf{ar{x_{4}}=3.2222}
mathbf{ar{x_{t}}=2.8889}

Tt: Total mean

ar{x_{1}}, ar{x_{2}},ar{x_{3}},ar{x_{4}}: Means :of :different :groups

As we know that ANOVA is used to study the variation between and within the groups.

F=rac{Difference:between:groups}{Difference:within:groups}

F=rac{MS_{b}}{MS_{w}}

where, MS_{w}=rac{SS_{w}}{df_{w}},:::MS_{b}=rac{SS_{b}}{df_{b}}

MS_{b}: mean square between groups

MS_{w}: mean square within groups.

df_{w}: degree of freedom for within group

df_{b}:degree of freedom for between group.

Calculation for mathbf{SS_{b}:}

mathbf{SS_{b}=Sigma N_{k}(ar{x_{k}}-ar{x_{t}})^{2}}

SS_{b}=N_{1}(ar{x_{1}}-ar{x_{t}})^{2}+N_{2}(ar{x_{2}}-ar{x_{t}})^{2}+N_{3}(ar{x_{3}}-ar{x_{t}})^{2}+N_{4}(ar{x_{4}}-ar{x_{t}})^{2}

SS_{b}=9(2.6667-2.8889)^{2}+9(2.2222-2.8889)^{2}+9(3.4444-2.8889)^{2}+9(3.2222-2.8889)^{2}

SS_{b}=0.4444+4.0004+2.7772+0.9998

SS, - 8.2218

Calculation for mathbf{SS_{T}}::

2

SS_{T=Sigma x_{1}^{2}+Sigma x_{2}^{2}+Sigma x_{3}^{2}+Sigma x_{4}^{2}-Nar{x}_{t}^{2}                    

N=4*9(no. of groups*no of dataset in each group)=36

SS_{T}=90+72+143+103-36(2.8889)^{2}

SST408-300.4468

SS_{T}=107.5532

SS_{w}=SS_{T}-SS_{b}

we have,

SS_{T}=107.5532,::SS_{b}=8.2218

SS_{w}=107.5532-8.2218

SS_{w}=99.3314

Calculation for degree of freedom df)

df_{w}=N-k=36-4=32

df_{b}=k-1=4-1=3

N: number of cases in dataset(i.e., 36=4*9 in our case)

k: number of groups(i.e., 4 in our case)

MS_{w}=rac{SS_{w}}{df_{w}} ,   Sb

since we have calculated all the desired value:

SSw 99.3314 dfu MS.-3 3.10410625 3.1041 32

MS,,,ー3.1041 ~. 3.10

MS, S 8.2218 dfb

MSb = 2.7406 ~ 2.74

So now we can calculate the value of mathbf{F:}

F=rac{MS_{b}}{MS_{w}}

F=rac{MS_{b}}{MS_{w}}=rac{2.74}{3.10}

Е 0.8839 0.88

F=0.88

The table for the calculated values:

Sources of variation SS df MS F
Between group 8.2218 approx 8.22 3 2.74 0.8839approx 0.88
Within group(error) 99.3314 approx 99.33 32 3.10
Total 107.5532 approx107.55 35